# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

### v1

@@ 5.2 Panda-gym
Task description. Our modified Panda-gym benchmark contains a robot arm control task using the Franka Emika Panda [Gallouédec et al., 2021], where the robot needs to move a cube to a target position. Unlike previous works, the robot can flexibly execute different skills (Push and Pick&Place) to solve tasks. In our experiments, we simultaneously modify multiple environmental features (cube mass and table friction) that characterise the transition dynamics. In some environmental settings,
Table 1: Comparison of success rate with baselines in Panda-gym (over 5 seeds).The blue background indicates that our method surpasses PEARL, TESAC and CCM. Bold text signifies the highest average return.
@@
TESAC 0.64$\pm$0.16 0.33$\pm$0.23 0.27$\pm$0.17
CCM 0.82$\pm$0.04 0.50$\pm$0.23 0.29$\pm$0.23
SaTESAC 0.92$\pm$0.05 0.55$\pm$0.25 0.40$\pm$0.35
SaCCM 0.94$\pm$0.02 0.58$\pm$0.27 0.46$\pm$0.36
only a specific skill can accomplish the task. For example, an agent needs to learn the Pick&Place skill to avoid high friction. Thus, the agent must determine when and which skill to execute.
Evaluating zero-shot generalisation. As shown in Table 5.1, SaTESAC and SaCCM achieve superior generalisation performance compared to PEARL, TESAC, and CCM, with a smaller sample space. In training, moderate test, and extreme test tasks, SaTESAC achieved up to 43.75%, 66.67%, and 48.15% higher success rates, respectively, compared to TESAC, while SaCCM improved success rates by 14.63%, 16.00%, and 58.62% compared to CCM.
Figure 6: (a) t-SNE visualisation of context embeddings in the Panda-gym environment. The clustered points in the yellow box correspond to the Push skill executed in scenarios with a large mass. Heatmap of (b) success rate, and probability of learned (c) Push skill and (d) Pick&Place skill of SaCCM. Under large mass scenarios, the probability of executing Push skill is higher than Pick&Place skill.
Skill analysis. We visualise the context embedding of SaCCM using t-SNE [Van der Maaten and Hinton, 2008]. We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table (see Appendix D for more details), and then employ Heatmaps [Waskom, 2021] to visualise skills executed by the agent. As shown in Figure 6, we found that when the cube mass is large (30 Kg and 10 Kg), the agent learned the Push skill (corresponding to the clustered points in the yellow bounding box in Figure 6(a). With smaller masses, the agent learned the Pick&Place skill. However, as shown in Figure 15 in Appendix F.1, CCM did not exhibit clear skill grouping. This indicates that SaMI extracts high-quality skill-related information from the trajectories and aids in embodying diverse skills. We provide additional visualisation based on both t-SNE and PCA [Jolliffe, 2002], and heatmap results in Appendix F.
5.3 MuJoCo
Task description. The modified MuJoCo benchmark [Mu et al., 2022, Lee et al., 2020] contains 6 typical robotic control environments based on the MuJoCo physics engine [Todorov et al., 2012]. Hopper, Half-cheetah, Ant, and SlimHumanoid are affected by multiple continuous environmental features (i.e., mass, damping) that characterise the transition dynamics. Both Crippled Ant and Crippled Half-cheetah are more difficult because they are also affected by a discrete environmental feature as one of the actuators in the legs is randomly crippled to change the transition dynamics.
Table 2: Comparison of average return with baselines in modified MuJoCo benchmark (over 5 seeds). The blue background indicates that our method surpasses PEARL, TESAC and CCM. Bold text signifies the highest average return.
@@ 5.3 MuJoCo
PEARL 182$\pm$73 96$\pm$21 88$\pm$31 2538$\pm$783 1028$\pm$445 908$\pm$185
TESAC 485$\pm$83 129$\pm$32 117$\pm$32 4512$\pm$737 1708$\pm$360 1550$\pm$274
CCM 814$\pm$23 146$\pm$13 158$\pm$31 4153$\pm$569 1417$\pm$710 995$\pm$188
SaTESAC 568$\pm$70 176$\pm$9 196$\pm$17 5380$\pm$917 2334$\pm$169 1848$\pm$341
SaCCM 719$\pm$82 215$\pm$6 219$\pm$9 6210$\pm$402 2656$\pm$231 2070$\pm$651
@@
PEARL 153$\pm$63 73$\pm$25 39$\pm$40 1802$\pm$773 530$\pm$270 346$\pm$102
TESAC 761$\pm$234 360$\pm$23 276$\pm$63 5817$\pm$3064 2843$\pm$1303 1074$\pm$446
CCM 1268$\pm$362 1078$\pm$492 548$\pm$177 6839$\pm$542 3053$\pm$347 1422$\pm$332
SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
SlimHumanoid Hopper
@@
PEARL 6947$\pm$3541 3697$\pm$2674 2018$\pm$907 934$\pm$242 874$\pm$366 799$\pm$298
TESAC 9095$\pm$894 7180$\pm$711 3709$\pm$35 1530$\pm$39 1476$\pm$9 1491$\pm$79
CCM 7297$\pm$790 5856$\pm$290 3501$\pm$877 1488$\pm$18 1420$\pm$46 1451$\pm$76
SaTESAC 10501$\pm$2102 8483$\pm$1502 6669$\pm$883 1500$\pm$24 1452$\pm$36 1449$\pm$15
SaCCM 9738$\pm$92 8389$\pm$231 5606$\pm$2027 1472$\pm$53 1456$\pm$13 1446$\pm$94
Results. Table 2 shows the average return of our method and baselines on both training and test tasks. Incorporating SaNCE into TESAC resulted in SaTESAC’s average return increasing by 19.41% in training tasks across all environments, 39.70% in moderate test tasks, and 65.40% in extreme test tasks. Similarly, adding SaNCE to CCM led to SaCCM achieving up to 31.69%, 44.83%, and 40.88% higher average returns in training tasks, moderate test tasks, and extreme test tasks, respectively. Overall, SaTESAC and SaCCM outperform baselines in both training and testing in most of the tasks, in which the Ant, Crippled Ant and Hopper tasks are exceptions. Comparing the visualisation of context embeddings for CCM and SACCM in the Ant, Crippled Ant, and Hopper environments (Appendix F.2), introducing SaNCE does not modify the grouping of context embeddings across different tasks. Notably in Hopper environment, all algorithms maintain consistent average returns when zero-shot generalising to unseen test tasks, with no clear clustering of context embeddings across different tasks (Table 22 in Appendix F.2). This indicates that a single skill may generalise across numerous tasks within the Hopper environment. Hence, we can draw an insight that SaMI is more suitable for environments that require multiple skills to adapt to different environmental features.
5.4 Analysis of the $\log$-$K$ curse in sample-limited scenarios
@@
In this section, we further analyse whether SaNCE can overcome the $\log$-$K$ curse. During the training phases, we sample the environmental features at the beginning of each episode. Therefore, throughout the training process, the context encoder needs to learn the context embedding distribution of multiple tasks. Because InfoNCE requires sampling negative samples across all tasks, and SaNCE samples negative samples from the current task, SaNCE’s negative sample space is $|c|$ times smaller than that of InfoNCE. For example, in the SlimHumanoid environment, where both mass and damping have five values, there can be a maximum of 25 different tasks, making the sampling space of InfoNCE potentially 25 times larger than that of SaNCE. According to Table 5.1 and 2, RL algorithms equipped with SaNCE can achieve better or comparable performance with a much smaller number of negative samples $K$ than InfoNCE. This indicates that SaNCE indeed plays a significant role in addressing the $\log$-$K$ curse, and the SaMI objective helps the contrastive context encoder extract information that is crucial for downstream RL tasks.
@@
Furthermore, the number of negative samples $K$ also relates to two hyperparameters: buffer size, which directly determines the size of the negative sample space; and contrastive batch size, which determines the number of samples used for training contrastive context encoder at each update step. Therefore, we conducted further analysis on these two hyperparameters. As shown in Figure 7, we find that reductions in buffer size and contrastive batch size do not significantly decrease the average return for SaCCM and SaTESAC, which exhibit state-of-the-art performance with small buffers and contrastive batch sizes. Note that the results in Table 2 correspond to a buffer size of 100,000 and a contrastive batch size of 12. Experimental results for all environments (Appendix E.2) further demonstrate that SaNCE is not highly sensitive to changes in $K$, and indeed shows great potential in overcoming the $\log$-$K$ curse.
@@ 6 Conclusion and future work
In this paper, we propose a Skill-aware Mutual Information (SaMI) to learn context embeddings for zero-shot generalisation in downstream RL tasks, and a Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI and overcome the $\log$-$K$ curse. RL algorithms equipped with SaMI have achieved state-of-the-art performance in the MuJoCo and Panda-gym benchmarks. Through visual analysis of context embeddings and heatmap analysis of learned skills in Panda-gym, we confirm that SaMI assists the RL agent in embodying diverse skills. These experimental observations indicate that the context encoder, learned by maximising SaMI, helps the RL agent generalise across various tasks by executing different skills. Notably, the optimisation process of SaNCE utilises a far smaller negative sample space than baselines. Coupled with further experimental analysis of buffer size and contrastive batch size on MuJoCo, we demonstrate that SaNCE helps overcome the $\log$-$K$ curse.
Figure 7: Changes of average return in the SlimHumanoid environment. (a) The impact of buffer size on RL control performance (i.e., TESAC, CCM, SaTESAC, SaCCM). (b) The impact of contrastive batch size on the contrastive context encoder (i.e., CCM, SaTESAC, SaCCM).
Our results indicate the importance of aligning the optimisation objectives of representation learning and downstream optimal control to enhance task performance and improve data efficiency.
Given that environmental features are often interdependent, such as a cube’s material correlating with friction and mass, SaMI does not introduce independence assumptions like DOMINO [Mu et al., 2022]. Therefore, future work will focus on verifying and enhancing SaMI’s potential in more complex tasks where environmental features are correlated. This will contribute to our ultimate goal: developing a generalist and versatile agent capable of working across multiple tasks and even real-world tasks in the near future.
@@ Appendix A Proof of Lemma 1
Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, A $K$-sample InfoNCE estimator is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$. InfoNCE is obtained by comparing positive pairs $(x,y_{1})$ and negative pairs $(x,y_{k})$, where $y_{k}\sim y_{2:K}$, as follows:
@@
Step 1. Let us prove the $K$-sample InfoNCE estimator is upper bounded by $\log K$. According to Mu et al. [2022], $\frac{f_{\psi}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}(x,y_{k})}=\frac{f_{\psi}(x,y_{1})}{f_{\psi}(x,y_{1})+\sum_{k=2}^{K}f_{\psi}(x,y_{k})}\leq 1$. So we have:
@@
Step 2. We have the $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K)$ which are rightmost and the leftmost side in Lemma 1 according to:
@@
[Poole et al., 2019] A $K$-sample estimator is an asymptotically tight lower bound to the MI, i.e.,
@@
Step 3. In this research, the context encoder $\psi$ in $f_{\psi}(x,y)$ is implemented using an RNN to approximate $\frac{p(y|x)}{p(y)}$ [Oord et al., 2019]. For most deep learning platforms, with a powerful learner for $\psi$ and a limited size $K$ such that $I(x;y)\geq\log K$, we can reasonably expect $I_{\text{InfoNCE}}\approx\log K$ after a few training epochs. Therefore, during training, when $K\ll+\infty$, we always have $I(x;y)\geq\log K$.
@@
Step 4. Let us prove that the $K$-sample InfoNCE bound is asymptotically tight. The specific choice of context encoder $\psi$ relates to the tightness of the $K$-sample NCE bound. InfoNCE [Oord et al., 2019] sets $f_{\psi}(x,y)\propto\frac{p(y|x)}{p(y)}$ to model a density ratio which preserves the MI between $x$ and $y$, where $\propto$ stands for ‘proportional to’ (i.e. up to a multiplicative constant). Let us plug in $f_{\psi}(x,y)=f_{\psi}^{*}(x,y)=\frac{p(y|x)}{p(y)}$ into InfoNCE, and we have
@@
Putting it together. Combining $I(x;y)\geq\log K$ with Proposition 1 and Equation 6, we have Lemma 1:
@@
Besides, according to Equation 7, with sample size $K\to+\infty$, $K$-sample InfoNCE bound is sharp and approaches to the true MI $I(x;y)$, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K\approx I(x;y)$.

### v3

@@ 5.2 Panda-gym
Task description. Our modified Panda-gym benchmark involves a robot arm control task using the Franka Emika Panda [Gallouédec et al., 2021], where the robot moves a cube to a target position. Unlike previous works, we simultaneously modify multiple environmental features (cube mass and table friction) that characterise the transition dynamics, and the robot can flexibly execute different skills (Push and Pick&Place) for different tasks. This environment requires high skill diversity; for instance, the agent must use Pick&Place in high-friction tasks and Push in high-mass tasks.
Table 1: Comparison of success rate $\pm$ standard deviation with baselines in Panda-gym (over 5 seeds). Bold text signifies the highest average return. * next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI. All significance claims based on paried t-tests with significance threshold of $p<0.05$.
@@
TESAC 0.50$\pm$0.22 0.31$\pm$0.20 0.22$\pm$0.21
CCM 0.80$\pm$0.19 0.49$\pm$0.23 0.29$\pm$0.28
SaTESAC 0.92$\pm$0.04 0.56$\pm$0.24 0.37$\pm$0.34
SaCCM 0.93$\pm$0.05 0.57$\pm$0.26 0.36$\pm$0.35
Results and skill analysis. As shown in Table 5.1, SaTESAC and SaCCM achieve superior generalisation performance compared to PEARL, TESAC, and CCM, with a smaller sample space. The t-test results in Table 5.1 show that SaMI significantly improves success rates across training, moderate, and extreme test sets at a 0.05 significance level. Video demos2 show that agents equipped with SaMI acquired multiple skills (Push, Pick&Place) to handle various tasks. When faced with an unknown task, the agents explore by attempting to lift the cube, infer the context, and adjust their skills accordingly within the episode. We visualised the context embeddings using UMAP [McInnes et al., 2020] (Figure 6(a) and Appendix F) and t-SNE [Van der Maaten and Hinton, 2008] (Appendix F), plotting the final step from 100 tests per task. Skills were identified through contact points between the end effector, cube, and table (see Appendix D for more details), and heatmaps [Waskom, 2021] were used to visualise executed skills. Figure 6 shows that SaCCM agents learned the Push skill for large cube masses (30 Kg, 10 Kg) and Pick&Place for smaller masses, while CCM showed no clear skill grouping (Figure 16 in Appendix F.1). Overall, SaMI incentivises agents to autonomously learn diverse skills, enhancing generalisation across a wider range of tasks. Specifically, through the cycle of effective exploration, context inference, and adaptation, diverse skills emerge solely from the data. Further visualisation results are in Appendix F.
Table 2: Comparison of average return $\pm$ standard deviation with baselines in modified MuJoCo benchmark (over 5 seeds). Bold number signifies the highest return. * next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI. All significance claims based on t-tests with significance threshold of $p<0.05$.
@@
PEARL 1682$\pm$73 996$\pm$21 888$\pm$31 1998$\pm$973 698$\pm$548 746$\pm$1092
TESAC 2139$\pm$90 1952$\pm$40 1048$\pm$124 3967$\pm$955 874$\pm$901 846$\pm$849
CCM 2361$\pm$114 2047$\pm$83 1527$\pm$301 3481$\pm$488 821$\pm$575 873$\pm$914
SaTESAC 2638$\pm$406 2379$\pm$528 2131$\pm$132 4328$\pm$1092 1143$\pm$664 1540$\pm$1094
SaCCM 2355$\pm$170 2310$\pm$314 2007$\pm$68 4478$\pm$1131 1007$\pm$568 1027$\pm$782
@@
PEARL 5153$\pm$581 3873$\pm$235 3802$\pm$409 5802$\pm$773 2190$\pm$970 1346$\pm$692
TESAC 6789$\pm$451 4705$\pm$279 4108$\pm$369 6298$\pm$2310 3173$\pm$1210 1159$\pm$338
CCM 6901$\pm$567 5179$\pm$902 4700$\pm$696 6955$\pm$788 3963$\pm$622 1325$\pm$269
SaTESAC 7314$\pm$545 5513$\pm$648 4940$\pm$531 7430$\pm$1026 4058$\pm$890 1780$\pm$102
SaCCM 7478$\pm$539 5717$\pm$488 5215$\pm$377 7154$\pm$965 3849$\pm$689 1926$\pm$218
SlimHumanoid HumanoidStandup
@@
PEARL 6947$\pm$3541 3697$\pm$2674 2018$\pm$907 95456$\pm$13445 63242$\pm$13546 64224$\pm$15467
TESAC 8437$\pm$1798 6989$\pm$1301 3760$\pm$308 158384$\pm$14455 153944$\pm$15046 74220$\pm$19980
CCM 7696$\pm$1907 5784$\pm$531 2887$\pm$1058 146480$\pm$33745 154601$\pm$16291 94991$\pm$15258
SaTESAC 10216$\pm$1620 7886$\pm$2203 6123$\pm$1403 178142$\pm$10081 168337$\pm$12123 133335$\pm$24607
SaCCM 9312$\pm$705 7430$\pm$1587 6473$\pm$2001 187930$\pm$19338 181033$\pm$14628 141750$\pm$27426
Hopper Crippled Hopper
PEARL 934$\pm$242 874$\pm$366 799$\pm$298 3091$\pm$298 2387$\pm$656 456$\pm$235
TESAC 1492$\pm$59 1499$\pm$35 1459$\pm$72 3575$\pm$192 3298$\pm$551 722$\pm$161
CCM 1484$\pm$54 1446$\pm$64 1452$\pm$58 3455$\pm$301 3409$\pm$239 1009$\pm$289
SaTESAC 1502$\pm$20 1453$\pm$39 1447$\pm$14 3391$\pm$84 3262$\pm$166 1839$\pm$130
SaCCM 1462$\pm$45 1462$\pm$14 1451$\pm$67 3449$\pm$103 3390$\pm$211 2059$\pm$221
Walker Crippled Walker
PEARL 7524$\pm$2455 3355$\pm$2555 1984$\pm$356 7899$\pm$2532 4377$\pm$2563 2965$\pm$1426
TESAC 7747$\pm$1772 4355$\pm$1530 2581$\pm$407 9908$\pm$1561 5929$\pm$1971 3041$\pm$912
CCM 8136$\pm$557 5476$\pm$803 2519$\pm$682 10317$\pm$1137 6233$\pm$1869 3098$\pm$821
SaTESAC 8675$\pm$752 5840$\pm$676 3632$\pm$404 10389$\pm$1031 8387$\pm$1291 4280$\pm$485
SaCCM 8361$\pm$586 5779$\pm$691 3481$\pm$332 10496$\pm$951 8235$\pm$1212 4824$\pm$839
5.3 MuJoCo
Task description. We extended the modified MuJoCo benchmark introduced in DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020]. It contains ten typical robotic control environments based on the MuJoCo physics engine [Todorov et al., 2012]. Hopper, Walker, Half-cheetah, Ant, HumanoidStandup, and SlimHumanoid are influenced by continuous environmental features (i.e., mass, damping) that affect transition dynamics. Crippled Ant, Crippled Hopper, Crippled Walker, and Crippled Half-cheetah are more challenging due to the addition of discrete environmental features (i.e., randomly crippled leg joints), requiring agents to master different skills (e.g., switching from running to crawling after a leg is crippled).
Results and skill analysis. Table 2 shows the average return of our method and baselines on training and test tasks. SaTESAC and SaCCM achieved higher returns in most tasks, except for Ant, Half-Cheetah, and Hopper, where only a single skill was needed. For instance, the Hopper robot learned to hop forward, adapting to different mass values. When environments become complex and require diverse skills for different tasks (Crippled Ant, Crippled Hopper, Crippled Half-Cheetah, SlimHumanoid, HumanoidStandup, and Crippled Walker), SaNCE brings significant improvements. For example, when the Ant robot has 3 or 4 legs available, it learns to roll to generalise across varying mass and damping. In more challenging zero-shot settings, when only 2 legs are available, the ant robot can no longer roll and it adapts by walking using its 2 healthy legs. This aligns with the results in Table 2, where SaMI significantly improved performance in extreme test sets. In summary, i) SaMI helps the RL agents to be versatile and embody multiple skills; ii) SaMI leads to increased returns during training and zero-shot generalisation, especially in environments that require different skills. Our video demos2 and visualisation results in Appendix F.2 show different skills in all environments.
5.4 Analysis of the $\boldsymbol{\log}$-$\boldsymbol{K}$ curse in sample-limited scenarios
@@
This section analyses whether SaNCE can overcome the $\log$-$K$ curse. During training, environmental features are sampled at the start of each episode, requiring the context encoder to learn the context embedding distribution across multiple tasks. Since InfoNCE samples negative examples from all tasks and SaNCE samples from the current task, SaNCE’s negative sample space is $|c|$ times smaller than InfoNCE’s. For instance, in the SlimHumanoid environment, where both mass and damping have five values, InfoNCE’s sampling space can be 25 times larger than SaNCE’s. As shown in Tables 5.1 and 2, RL algorithms using SaNCE achieve better or comparable performance with significantly fewer negative samples ($K$) than InfoNCE. This suggests SaNCE effectively addresses the $\log$-$K$ curse, and the SaMI objective helps the contrastive context encoder extract critical information for downstream RL tasks.
@@
The number of negative samples $K$ is influenced by two hyperparameters: buffer size, which determines the negative sample space, and contrastive batch size, which controls the number of samples used to train the contrastive context encoder per update. We analysed these hyperparameters further, and as shown in Figure 7, reductions in buffer and contrastive batch size do not significantly impact the average return for SaCCM and SaTESAC, which maintain state-of-the-art performance with small buffers and batch sizes. The results in Table 2 correspond to a buffer size of 100,000 and a batch size of 12. Results across all environments (refer to Appendix E.2) show that SaNCE exhibits low sensitivity to $K$, highlighting its potential to overcome the $\log$-$K$ curse.
Figure 7: Effect of (a) buffer size (TESAC, CCM, SaTESAC, SaCCM) and (b) contrastive batch size (CCM, SaTESAC, SaCCM) in the SlimHumanoid environment.
@@ 6 Conclusion and future work
Zero-shot generalisation has been a longstanding challenge concerning Meta-RL agents, with skill diversity and sample efficiency being key to generalising to previously unseen environments. In this paper, we proposed Skill-aware Mutual Information (SaMI) to learn context embeddings for zero-shot generalisation in downstream RL tasks, and Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI and overcome the $\log$-$K$ curse, along with a practical skill-aware trajectory sampling strategy. Experimental results showed that RL algorithms equipped with SaMI achieved state-of-the-art performance in MuJoCo and Panda-gym benchmarks, particularly in zero-shot generalisation within more complex environments. During the zero-shot generalisation, when faced with an unseen task, SaMI assists agents in exploring effectively, inferring context, and rapidly adapting their skills within the current episode. SaNCE’s optimisation uses a significantly smaller negative sample space than baselines, and our analysis on buffer and contrastive batch sizes demonstrated its effectiveness in addressing the $\log$-$K$ curse.
Given that environmental features are often interdependent, such as a cube’s material correlating with friction and mass, SaMI does not introduce independence assumptions like DOMINO [Mu et al., 2022]. Therefore, future work will focus on verifying and enhancing SaMI’s potential in more complex tasks where environmental features are correlated [Dunion et al., 2023a]. This will contribute to our ultimate goal: developing a generalist and versatile agent capable of working across multiple tasks and even real-world tasks in the near future.
@@ Appendix A Proof of Lemma 1
Given a query $x$ and a set $Y=\{y_{1},\dots,y_{K}\}$ of $K$ random samples, containing one positive sample $y_{1}$ and $K-1$ negative samples drawn from the distribution $p(y)$, a $K$-sample InfoNCE estimator is obtained by comparing pairs sampled from the joint distribution $(x,y_{1})\sim p(x,y)$ with pairs $(x,y_{k})$, constructed using the set of negative examples $y_{2:K}$. InfoNCE compares the positive pairs $(x,y_{1})$ with the negative pairs $(x,y_{k})$, where $y_{k}\sim y_{2:K}$, as follows:
@@
Step 1. Let us prove that the $K$-sample InfoNCE estimator is upper-bounded by $\log K$. According to Mu et al. [2022], $\frac{f_{\psi}(x,y_{1})}{\sum_{k=1}^{K}f_{\psi}(x,y_{k})}=\frac{f_{\psi}(x,y_{1})}{f_{\psi}(x,y_{1})+\sum_{k=2}^{K}f_{\psi}(x,y_{k})}\leq 1$. So we have:
@@
Step 2. We have the $I(x;y)\geq I_{\text{InfoNCE}}(x;y|\psi,K)$ according to:
@@
[Poole et al., 2019] A $K$-sample estimator is an asymptotically tight lower bound on the MI, i.e.,
@@
Step 3. In this research, the context encoder $\psi$ in $f_{\psi}(x,y)$ is implemented using an RNN to approximate $\frac{p(y|x)}{p(y)}$ [Oord et al., 2019]. With a sufficiently powerful deep learning model for $\psi$ and a finite sample size $K$, such that $I(x;y)\geq\log K$, we can reasonably expect that $I_{\text{InfoNCE}}\approx\log K$ after a few training epochs. Therefore, during training, when $K\ll+\infty$, we always have $I(x;y)\geq\log K$.
@@
Step 4. Let us prove that the $K$-sample InfoNCE bound is asymptotically tight. The specific choice of the context encoder $\psi$ influences the tightness of the $K$-sample NCE bound. InfoNCE [Oord et al., 2019] sets $f_{\psi}(x,y)\propto\frac{p(y|x)}{p(y)}$ to model a density ratio that preserves the MI between $x$ and $y$, where $\propto$ stands for ’proportional to’ (i.e., up to a multiplicative constant). Substituting $f_{\psi}(x,y)=f_{\psi}^{*}(x,y)=\frac{p(y|x)}{p(y)}$ into InfoNCE, we obtain:
@@
Putting it together. Combining $I(x;y)\geq\log K$ with Proposition 1 and Equation (6), we have Lemma 1:
@@
Moreover, according to Equation (7), as the sample size $K\to+\infty$, the $K$-sample InfoNCE bound becomes sharp and approaches the true MI $I(x;y)$, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K\approx I(x;y)$.